OpenStudy (anonymous):

find the complex zero of the polynomial function. Write f in factored form. f(x)=x^4+10x^2+9 f(x)=_________________________

OpenStudy (anonymous):

it cannot be factored out

OpenStudy (anonymous):

really but it says that i should write it in a factored form. and how would you find the complex zero?

OpenStudy (anonymous):

so do I just but it in the quadratic form when i get the x^2

OpenStudy (campbell_st):

it can be factored \[(x^2 + 9)(x^2 + 1) = 0 \] so the complex roots occur when \[x^2 + 9 = 0.....and.....x^2 + 1 = 0\] hopefullt this helps you find them

OpenStudy (campbell_st):

so do you know how to now find the complex roots..?

OpenStudy (anonymous):

so it would be either x= -sqrt of 9 and - sqrt of 1

OpenStudy (campbell_st):

well not quite here is the start using i^2 = -1 \[x^2 = - 9 ... x = \sqrt{-9}.... x = \sqrt{9i^2}\]

OpenStudy (campbell_st):

so one pair of complex roots is \[x = \pm 3i\] you just need to find the other 2 complex roots.

OpenStudy (anonymous):

oohhhhh okay but it would be +-3i and the other one would be +-1i so it would just be +-i

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

so I got the answer for the complex zeros but how would it be in factored form??

OpenStudy (campbell_st):

well then you will have 4 binomials (x - 3i)(x+3i)... I'll let you get the other 2 binomials.

OpenStudy (anonymous):

okay so it would be (x-i)(x+i) and so do I but them all together like (x-3i)(x+3i)(x-i)(x+i)

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

thank you for taking me through the steps!!

OpenStudy (campbell_st):

glad to help